40m QIL Cryo_Lab CTN SUS_Lab CAML OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 289 of 350  Not logged in ELOG logo
ID Date Author Type Category Subject
  3082   Wed Jun 16 18:14:13 2010 AidanUpdateGeneralGlass cover from overhead light smashed on PSL table

I was giving a tour of the 40m yesterday. We were looking at the PSL table. About 30 seconds after I turned the lights on a glass cover from one of the lights (NW corner) popped out of its holder and smashed on the table.

I've cleaned up all the broken glass I could see but there may be some small shards there. Please use caution in that area.

  3081   Wed Jun 16 18:12:16 2010 nancyConfigurationComputers40MARS

i added my laptop's mac address to teh martian at port 13 today.

 

  3080   Wed Jun 16 11:31:19 2010 josephbSummaryComputersRemoved scaling fonts from medm on Allegra

Because it was driving me crazy while working on the new medm screens for the simulated plant, I went and removed the aliased font entries in /usr/share/X11/fonts/misc/fonts.alias that are associated with medm.  Specifically I removed the lines  starting with widgetDM_.  I made a backup in the same directory called fonts.alias.bak with the old lines.

Medm now behaves the same on op440m, rosalba, and allegra - i.e. it can't find the widgetDM_ scalable fonts and defaults to a legible fixed font.

  3079   Tue Jun 15 21:28:44 2010 albertoUpdateLocking40m Upgrade Optickle Model

Quote:

Quote:

 

 In my calculation of the digital filters of the optical transfer functions the carrier light is resonant in coupled cavities and the sidebands are resonant in recycling cavities (provided that macroscopic lengths are chosen correctly which I assumed).

Carrier and SB (f2) shouldn't be resonant at the same time in the SRC-arms coupled cavity. No additional filtering of the GW signal is wanted.

The SRC macroscopic length is chosen to be = c / f2 - rather than = [ (n+1/2) c / (2*f2) ] - accordingly to that purpose.

I calculated the frequency of the double cavity pole for the 40m SRC-arm coupled cavity.

w_cc = (1 + r_srm)/(1- r_srm) * w_c

where w_c is the arm cavity pole angular frequency [w_c = w_fsr * (1-r_itm * r_etm)/sqrt(r_itm*r_etm) ]

I found the pole at about 160KHz. This number coincides with what I got earlier with my optickle model configured and tuned as I said in my previous entry. See attachments for plots of transfer functions with 0 and 10pm DARM offsets, respectively.

I think  the resonance at about 20 Hz that you can see in the case with non-zero DARM offset, is due to radiation pressure. Koji suggested that I could check the hypothesis by changing either the mirrors' masses or the input power to the interferometer. When I did it frequency and qualty factor of the resonance changed, as you would expect for a radiation pressure effect.

Independently, Jan also calculated the pole frequency of the transfer function DARM / ASQ2 as we would expect it for the SRC-coupled cavity. He also found the pole at about 160KHz. I'm attaching the plot with the transfer function he calculated.
He also said that the little bump at the pole frequency is OK considering that our signal recycling cavity is not much shorter than the arms.

This gave me more confidence about my optickle model of the 40m. This is quite comforting since I used that model other times in the past to calculate several things (i.e. effects of higher unwanted harmonics from the oscillator, or, recently, the power at the ports due to the SB resonating in the arms).

I don't know anymore what Valera said that wasn't right.
Also, as he said, he set it for the carrier to be resonant in the SRC-arms couple cavity. But that is not our case.
  3078   Tue Jun 15 19:24:25 2010 ranaUpdateIOOMode Profile after Mode Matching Telescope

Quote:

We obtained a good mode match overlap of 99.0% for the new IOO.

And if we move the position of MMT2 by another 10 cm away from MMT1, we will have 99.6% overlap.

 That's hot stuff.

  3077   Tue Jun 15 16:28:32 2010 kiwamuUpdateIOOMode Profile after Mode Matching Telescope

We obtained a good mode match overlap of 99.0% for the new IOO.

And if we move the position of MMT2 by another 10 cm away from MMT1, we will have 99.6% overlap. 

Yesterday Jenne and I put MMT2 on the OMC table. MMT2 was carefully put by measuring the distance between MMT1 and MMT2.

The position looked almost the same as that drawn on the CAD design.

After putting it we measured the profile after the MMT.

 

The attached figure shows the computed mode overlap according to the fitting result while changing the position of MMT2 in a program code.

The x-axis is the position of MMT2, the current position is set to be zero. The y-axis is the mode match overlap.

Right now the overlap is 99.0% successfully, but this is not an optimum point because the maximum overlap can be achieved at x=100 mm in the plot.

It means we can have 99.6% by moving the position of MMT2 by another 10 cm. This corresponds to an expansion of the MMT length.

If this expansion is difficult due to the narrow available space in the chamber, maybe staying of MMT2 at the current position is fine.

  3076   Mon Jun 14 22:16:08 2010 JenneUpdateIOOMode scan after Mode Matching Telescope

[Jenne, Kiwamu]

We measured the mode after the Mode Matching Telescope. 

 

---- fitted parameters ----

       w0_h =  2.85 +/- 0.0115 mm

       w0_v =  2.835 +/- 0.00600 mm

       z0_h = 5.4 +/- 0.447  m

       z0_v  = 6.9 +/- 0.305  m

  3075   Mon Jun 14 07:57:07 2010 albertoUpdateLocking40m Upgrade Optickle Model

Quote:

 

 In my calculation of the digital filters of the optical transfer functions the carrier light is resonant in coupled cavities and the sidebands are resonant in recycling cavities (provided that macroscopic lengths are chosen correctly which I assumed).

Carrier and SB (f2) shouldn't be resonant at the same time in the SRC-arms coupled cavity. No additional filtering of the GW signal is wanted.

The SRC macroscopic length is chosen to be = c / f2 - rather than = [ (n+1/2) c / (2*f2) ] - accordingly to that purpose.

  3074   Sun Jun 13 08:28:44 2010 valeraUpdateLocking40m Upgrade Optickle Model

 In my calculation of the digital filters of the optical transfer functions the carrier light is resonant in coupled cavities and the sidebands are resonant in recycling cavities (provided that macroscopic lengths are chosen correctly which I assumed).

  3073   Sat Jun 12 19:43:19 2010 AlbertoUpdateWIKI-40M UpdateIFO modeling Wiki Page updated

Today I started writing the IFO modeling wiki page.

The idea is to make it a reference place where to share our modeling tools for the 40m.

  3072   Sat Jun 12 19:41:04 2010 AlbertoUpdateLocking40m Upgrade Optickle Model

I wrote down the settings according to which I tuned the optickle model of the 40m Upgrade.

Basically I set it so that:

  1. PRC alone anti-resonant for the carrier and resonant for both sidebands
  2. SRC alone resonant for the carrier and resonant for the f2 sideband

In this way when the carrier becomes resonant in the arms we have:

  1. carrier resonant in PRC and anti-resonant in SRC
  2. f1 resonant in PRC and non resonant in SRC
  3. f2 resonant in SRC

The DARM offset for DC readout is optional, and doesn't change those conditions.

I also plotted the carrier and the sideband's circulating power for both recycling cavities.

I'm attaching a file containing more detailed explanations of what I said above. It also contains the plots of field powers, and transfer functions from DARM to the dark port. I think they don't look quite right. There seems to be something wrong.

Valera thought of fixing the problem, removing the 180 degree offset on the SRM, which is what makes the sideband rather than the carrier resonant in SRC. In his model the carrier becomes resonant and the sideband anti-resonant. I don't think that is correct.

The resonant-carrier case is also included in the attachment (the plots with SRMoff=0 deg). In the plots the DARM offset is always zero.

I'm not sure why the settings are not producing the expected transfer functions.

  3071   Sat Jun 12 18:03:00 2010 sharmilaUpdateelogTemperature Controller

Kiwamu and I setup a serial port terminal for receiving data from TC200 via a RS-232 USB interface. It was done using a Python code. Some command definitions need to be done to get the output from TC-200.

  3070   Fri Jun 11 22:09:58 2010 valeraHowToCDSfoton

 It appears that foton does not like the unstable poles, which we need to model the transfer functions.

But one can try to load the filters into the front end by generating the filter file e.g.:

#
# MODULES DARM_ASDC
 
#
################################################################################
### DARM_ASDC                                                                   ###
################################################################################
# SAMPLING DARM_ASDC  16384
# DESIGN   DARM_ASDC  
### ####
DARM_ASDC  0 21 6  0  0 darm 1014223594.005454063416 -1.95554205062071  0.94952075557861 0.06176931505784 -0.93823068494216
                         -2.05077577179611   1.05077843532639  -2.05854170261687  1.05854477394411
                         -1.85353637553024   0.86042048250739  -1.99996540107622  0.99996542454814 
                         -1.93464836371852   0.94008893626414  -1.89722830906561  0.90024221050918
                         -2.04422931770060   1.04652211283968  -2.01120153956052  1.01152717233685 
                         -1.99996545575365   0.99996548582538  -1.99996545573320  0.99996548582538

 

 

 

Unfortunately if you open and later save this file with foton it will strip the lhp poles.

  3069   Fri Jun 11 15:04:25 2010 josephbUpdateCDSMulti-filter matrix medm screens finished and script for copying filters from SOS file

I've finished the MEDM portion of the RCG FiltMuxMatrix part.  Now it generates an appropriate medm screen for the matrix, with links to all the filter banks.  The filter bank .adl files are also generated, and placed in a sub directory with the name of the filter matrix as the name of the sub directory.

The input is the first number and the output is the second number.  This particular matrix has 5 inputs (0 through 4) and 15 outputs (0 through 14). Unfortunately, the filter names can't be longer than 24 characters, which forced me to use numbers instead of actual part names for the input and output.

 

The key to the numbers is:

Inputs:

DARM 0
MICH 1
PRC 2
SRC 3
CARM 4

Outputs:

AS_DC 0
AS11_I 1
AS11_Q 2
AS55_I 3
AS55_Q 4
POP_DC 5
POP11_I 6
POP11_Q 7
POP55_I 8
POP55_Q 9
REFL_DC 10
REFL11_I 11
REFL11_Q 12
REFL55_I 13
REFL55_Q 14

To get this working required modifications to the feCodeGen.pl and the creation of mkfiltmatrix.pl (which was based off of mkmatrix.pl).  These are located in /cvs/cds/caltech/cds/advLigoRTS/src/epics/util/

In related news, I asked Valera if he could load the simulated plant filters he had generated, and after several tries, his answer was no.  He says it has the same format as those filter they pass to the feed forward banks down in Livingston, so he's not sure why they won't work.

I tested my script, FillFotonFilterMatrix.py, on some simple second order section filters (like gain of 1, with b1 = 1.01, b2 = 0.02, a1 = 1.03, a2 = 1.04), and it populated the foton filter file correctly, and was parsed fine by Foton itself.  So I'm going to claim the script is done and its the fault of the filters we're trying to load.  This script is now living in /cvs/cds/caltech/chans/ , along with a name file called lsp_dof2pd_mtrx.txt which tells the script that DARM is 0, CARM is 1, etc.  To run it, you also need a SOS.txt file with the filters to load, similar to the one Valera posted here, but preferably loadable.

I also updated my progess on the wiki, here.

  3068   Fri Jun 11 14:31:04 2010 kiwamuUpdateIOOmode matching of new IOO

We decided not to care about the mode after MMT1.

So far Koji, Alberto and I have measured the beam profile after MMT1,

but we are going to stop this measurement and go ahead to the next step i.e. putting MMT2

There are two reasons why we don't care about the profile after MMT1:

     (1) it is difficult to fit the measured data

     (2) The position of MMT1 is not critical for the mode matching to the IFO.

The details are below.

 


(1) difficulty in fitting the data

The precision of each measured point looked good enough, but the fitting result varies every measurement.

The below shows the data and their fitted curves. 

 profile_MMT.png

In the label, "h" and "v" stand for "horizontal" and "vertical" respectively.

The solid curves represent the fitting results, varying by each measurement.

In order to increase the reliability of the fitting, we had to take some more data at further distance.

But we couldn't do it, because the beam radius already becomes 3 mm even at 2 m away from MMT1 and at this point it starts to be clipped on the aperture of the beam scan.

Thus it is difficult to increase the reliability of the fitting. 

Once if we put MMT2 the beam should have a long Rayleigh range, it means we can measure the profile at further distance, and the fitting must be more reliable.

 


(2) positioning of MMTs

Actually the position of MMT1 is not so critical for the mode matching. 

The most important point is the separation distance of MMT1 and MMT2.

As written in Jenne's document, if we slide the positions of MMT1 and MMT2 while keeping their appropriate separation distance, the mode match overlap still stays above 99%

This is because the beam coming from MC3 is almost collimated (ZR~8m), so the position of MMTs doesn't so matter. 

To confirm it for the real case, I also computed the mode overlap while sliding the position of MMTs by using real data. The below is the computed result.

dist_MMO.png

It is computed by using the measured profile after MC3 (see the past entry).

The overlap still stay above 99% when the distance from MC to MMT is between 1300 and 3000mm.

This result suggests to us putting MMT1 as we like.  

  3066   Fri Jun 11 13:32:28 2010 KojiUpdateGreen Lockingend PDH with thermal feedback

GJ!

Quote:

A thermal feedback was installed to the end PDH locking and it works well. There are no saturations 

As I said the feedback signal was sometimes saturated at the sum-amp because the drive signal going to the laser PZT was large at low frequency (below 1Hz).

So I made a passive low pass filter which filters the signal controlling the temperature of the laser crystal, and put it before the temperature drive input.

Now the amount of the feedback signal got reduced when it is locked, and there are no saturations. It's very good.

 

  3065   Fri Jun 11 11:54:42 2010 kiwamuUpdateGreen Lockingend PDH with thermal feedback

A thermal feedback was installed to the end PDH locking and it works well. There are no saturations 

As I said the feedback signal was sometimes saturated at the sum-amp because the drive signal going to the laser PZT was large at low frequency (below 1Hz).

So I made a passive low pass filter which filters the signal controlling the temperature of the laser crystal, and put it before the temperature drive input.

Now the amount of the feedback signal got reduced when it is locked, and there are no saturations. It's very good.


(thermal property of the crystal) 

According to the specification sheet for the 1W Innolight, the thermal properties of the crystal are:

  Response coefficient : 3GHz/K 

  Temperature control coefficient : 1K/V

  Thermal response bandwidth: 1Hz 

 

(filter circuit and actuator response)

In order to feedback the signal blow 1Hz, a low pass fiter is needed. 

The attachment:1 shows the filter circuit I made.

Since I found that the drive input had an input impedance of 100kOhm, so I put relatively big resistors to have a moderate gain.

The expected actuator responses are also attached.

The blue curve represents the response of the PZT, the green is the thermal response including the low pass filter and the red curve is the total response composed of both the responses.

I assume that the PZT response is 1MHz/V according to Mott's measurement.

Also I assume that the thermal response intrinsically has two poles at 1Hz according to the specification listed above.

In the total response, there is a little gain reduction around 2Hz due to the cancelation of each other, but it still looks okay.

 

  3064   Thu Jun 10 11:10:21 2010 AlbertoUpdateGeneralLaTeXlabs

Quote:

I could not dare to share my google doc with this site...

Quote:

Quote:

BTW, latex launched this new thing for writing pdfs. doesnot require any installations.  check  http://docs.latexlab.org

 so cool!

 

Just in case,  granted access to Google docs can be revoked any time from here:

https://www.google.com/accounts/IssuedAuthSubTokens

  3063   Thu Jun 10 10:58:02 2010 KojiUpdateGeneralLaTeXlabs

I could not dare to share my google doc with this site...

Quote:

Quote:

BTW, latex launched this new thing for writing pdfs. doesnot require any installations.  check  http://docs.latexlab.org

 so cool!

 

  3062   Thu Jun 10 07:53:14 2010 AlbertoUpdatePEMLaTeXlabs

Quote:

BTW, latex launched this new thing for writing pdfs. doesnot require any installations.  check  http://docs.latexlab.org

 so cool!

  3061   Wed Jun 9 21:05:44 2010 ranaSummaryComputersop540m is not to be used

This is a reminder (mainly for Steve, who somehow doesn't believe these things) that op540m is not to be used for your general pleasure.

No web, no dataviewer, no DTT. Using these things often makes the graphical X-Windows crash. I have had to restart the StripTool, our seismic BLRMS and our Alarms many times because someone uses op540m, makes it crash, and then does not restart the processes.

Stop breaking op540m, Steve!

  3060   Wed Jun 9 19:47:08 2010 nancyUpdatePEMlead balls on concrete

Quote:

Quote:

Valera and I put the 2 Guralps and the Ranger onto the big granite slab and then put the new big yellow foam box on top of it.

There is a problem with the setup. I believe that the lead balls under the slab are not sitting right. We need to cut out the tile so the thing sits directly on some steel inserts.

You can see from the dataviewer trend that the horizontal directions got a lot noisier as soon as we put the things on the slab.

 The tiles were cut out in 1.5" ID circle to insure that the 7/16" OD lead balls would not touch the tiles on Wednesday, May 26, 2010

Granite surface plate specifications: grade B, 18" x 24" x 3" , 139 lbs

These balls and granite plate were removed by  Rana in entry log #3018 at 5-31-2010

 I tried to calculate the frequency of resonance using Rayleigh's method.  approximated the geometry of lead to be that of a perfect cylinder, and  the deformation in the lead by the deflection in a cantilever under  a shear strain.

this rough calculation gives an answer of 170Hz and depends on the dimensions of each lead, number of leads, and mass of the granite. But the flaw pointed out is that this calculation doesnot depend on the dimension of the granite slab, nor on the exact placing of the lead spheres with respect toteh COM of the slab.

I will put up the calculations details later, and also try to do a FEM analysis of the problem.

 

BTW, latex launched this new thing for writing pdfs. doesnot require any installations.  check  http://docs.latexlab.org

  3059   Wed Jun 9 11:13:11 2010 kiwamuUpdateGreen Lockinglock with PDH box

A progress on the end PDH locking :

by using a modified PDH box the green laser on the X-end station is locked to the arm cavity.

So far the end PDH locking had been achieved by using SR560s, but it was not sophisticated filter.

To have a sophisticated filter and make the control loop more stable, the PDH box labeled "#G1" was installed instead of the SR560s.

After the installation the loop looks more stable than the before.

Some details about the modification of the PDH box will be posted later.

 

Although, sometimes the loop was unlocked because the sum-amp (still SR560) which mixes the modulation and the feedback signal going to the NPRO PZT was saturated sometimes.

Thus as we expected a temperature control for the laser crystal is definitely needed in order to reduce such big low frequency drive signal to the PZT.

  3058   Wed Jun 9 10:31:13 2010 steveUpdateSAFETYsafety training

Nancy and Sharmila received introductory early bird surf safety training for the 40m lab.

  3057   Tue Jun 8 20:52:25 2010 josephbUpdatePEMDAQ up (for the moment)

As a test, I did a remote reboot of both Megatron and c1iscex, to make sure there was no code running that might interfere with the dataviewer.  Megatron is behind a firewall, so I don't see how it could be interfering with the frame builder.  c1iscex was only running a test module from earlier today when I was testing the multi-filter matrix part.  No daqd or similar processes were running on this machine either, but it is not behind a firewall at the moment. 

Neither of these seemed to affect the lack of past data.  I note the error message from dataviewer was "read(); errno=9".

Going to the frame builder machine, I ran dmesg.  I get some disturbing messages from May 26th and June 7th. There are 6-7 of these pairs of lines for each of these days, spread over the course of about 30 minutes.

Jun 7 14:05:09 fb ufs: [ID 213553 kern.notice] NOTICE: realloccg /: file system full

Jun 7 14:11:14 fb last message repeated 19 times

There's also one:

Jun 7 13:35:14 fb syslogd: /usr/controls/main_daqd.log: No space left on device

I went to /usr/controls/ and looked at the file.  I couldn't read it with less, it errored with Value too large for defined data type.  Turns out the file was 2.3 G.  And had not been updated since June 7th.  There were also a bunch of core dump files from May 25th, and a few more recent.  However the ones from May 25th were somewhat large, half a gig each or so.  I decided to delete the main_daqd.log file as well as the core files.

This seems to have fixed the data history for the moment (at least with one 16k channel I tested quickly). However, I'm now investigating why that log file seems to have filled up, and see if we can prevent this in the future.

Quote:

As before, I am unable to get data from the past. With DTT on Allegra I got data from now, but its unavailable from 1 hour ago. Same problem using mDV on mafalda. I blame Joe again - or the military industrial complex.

 

Quote:

Quote:

 Although trends are available, I am unable to get any full data from in the past (using DTT or DV). I started the FB's daqd process a few times, but no luck. 

I blame Joe's SimPlant monkeying from earlier today for lack of a better candidate. I checked and the frames are actually on the FB disk, so its something else.

 I tried running dataviewer and dtt this morning.  Dataviewer seemed to be working.  I was able to get trends, full data on a 2k channel (seismic channels) and full data on a 16k channel (C1:PEM-AUDIO_MIC1)  This was tried for a period 24 hours a go for a 10 minute stretch.

I also tried dtt and was able to get 2k and 16k channel data, for example C1:PEM-AUDIO_MIC1.  Was this problem fixed by someone last night or did time somehow fix it?

 

  3056   Tue Jun 8 18:39:36 2010 ranaUpdatePEMDAQ down

As before, I am unable to get data from the past. With DTT on Allegra I got data from now, but its unavailable from 1 hour ago. Same problem using mDV on mafalda. I blame Joe again - or the military industrial complex.

 

Quote:

Quote:

 Although trends are available, I am unable to get any full data from in the past (using DTT or DV). I started the FB's daqd process a few times, but no luck. 

I blame Joe's SimPlant monkeying from earlier today for lack of a better candidate. I checked and the frames are actually on the FB disk, so its something else.

 I tried running dataviewer and dtt this morning.  Dataviewer seemed to be working.  I was able to get trends, full data on a 2k channel (seismic channels) and full data on a 16k channel (C1:PEM-AUDIO_MIC1)  This was tried for a period 24 hours a go for a 10 minute stretch.

I also tried dtt and was able to get 2k and 16k channel data, for example C1:PEM-AUDIO_MIC1.  Was this problem fixed by someone last night or did time somehow fix it?

  3055   Tue Jun 8 15:58:25 2010 josephb, alexUpdateCDSNew multi-filter matrix part added to RCG (at the 40m at least)

A new webview of the LSP model is available at:

https://nodus.ligo.caltech.edu:30889/FE/lsp_slwebview_files/

This model include a couple example noise generators as well as the new Matrix of Filter banks (5 inputs x 15 outputs = 75 Filters!).  The attached png shows where these parts can be found in the CDS_PARTS library.  I'm still working on the automatic generation of the matrix and filter bank medm screens for this part.  The plan is to have a matrix screen similar to current ones, except that the value entry points to the gain setting of the associated filter.  In addition, underneath each value, there will be a link to the full filter bank screen.  Ideally, I'd like to have the filter adl files located in a sub-directory of the system, to keep clutter down.

I've cut and past the new Foton file generated by the LSP model below.  The first number following the MTRX is the input the filter is taking data from and the second number is the output its pushing data to.  This means for the script parsing Valera's transfer functions, I need to input which channel corresponds to which number, such as DARM = 0, MICH = 1, etc.  So the next step is to write this script and populate the filter banks in this file.

# FILTERS FOR ONLINE SYSTEM
#
# Computer generated file: DO NOT EDIT
#
# MODULES DOF2PD_AS11I DOF2PD_AS11Q DOF2PD_AS55I DOF2PD_AS55Q
# MODULES DOF2PD_ASDC DOF2PD_POP11I DOF2PD_POP11Q DOF2PD_POP55I
# MODULES DOF2PD_POP55Q DOF2PD_POPDC DOF2PD_REFL11I DOF2PD_REFL11Q
# MODULES DOF2PD_REFL55I DOF2PD_REFL55Q DOF2PD_REFLDC Mirror2DOF_f2x1
# MODULES Mirror2DOF_f2x2 Mirror2DOF_f2x3 Mirror2DOF_f2x4 Mirror2DOF_f2x5
# MODULES Mirror2DOF_f2x6 Mirror2DOF_f2x7 DOF2PD_MTRX_0_0 DOF2PD_MTRX_0_1
# MODULES DOF2PD_MTRX_0_2 DOF2PD_MTRX_0_3 DOF2PD_MTRX_0_4 DOF2PD_MTRX_0_5
# MODULES DOF2PD_MTRX_0_6 DOF2PD_MTRX_0_7 DOF2PD_MTRX_0_8 DOF2PD_MTRX_0_9
# MODULES DOF2PD_MTRX_0_10 DOF2PD_MTRX_0_11 DOF2PD_MTRX_0_12 DOF2PD_MTRX_0_13
# MODULES DOF2PD_MTRX_0_14 DOF2PD_MTRX_1_0 DOF2PD_MTRX_1_1 DOF2PD_MTRX_1_2
# MODULES DOF2PD_MTRX_1_3 DOF2PD_MTRX_1_4 DOF2PD_MTRX_1_5 DOF2PD_MTRX_1_6
# MODULES DOF2PD_MTRX_1_7 DOF2PD_MTRX_1_8 DOF2PD_MTRX_1_9 DOF2PD_MTRX_1_10
# MODULES DOF2PD_MTRX_1_11 DOF2PD_MTRX_1_12 DOF2PD_MTRX_1_13 DOF2PD_MTRX_1_14
# MODULES DOF2PD_MTRX_2_0 DOF2PD_MTRX_2_1 DOF2PD_MTRX_2_2 DOF2PD_MTRX_2_3
# MODULES DOF2PD_MTRX_2_4 DOF2PD_MTRX_2_5 DOF2PD_MTRX_2_6 DOF2PD_MTRX_2_7
# MODULES DOF2PD_MTRX_2_8 DOF2PD_MTRX_2_9 DOF2PD_MTRX_2_10 DOF2PD_MTRX_2_11
# MODULES DOF2PD_MTRX_2_12 DOF2PD_MTRX_2_13 DOF2PD_MTRX_2_14 DOF2PD_MTRX_3_0
# MODULES DOF2PD_MTRX_3_1 DOF2PD_MTRX_3_2 DOF2PD_MTRX_3_3 DOF2PD_MTRX_3_4
# MODULES DOF2PD_MTRX_3_5 DOF2PD_MTRX_3_6 DOF2PD_MTRX_3_7 DOF2PD_MTRX_3_8
# MODULES DOF2PD_MTRX_3_9 DOF2PD_MTRX_3_10 DOF2PD_MTRX_3_11 DOF2PD_MTRX_3_12
# MODULES DOF2PD_MTRX_3_13 DOF2PD_MTRX_3_14 DOF2PD_MTRX_4_0 DOF2PD_MTRX_4_1
# MODULES DOF2PD_MTRX_4_2 DOF2PD_MTRX_4_3 DOF2PD_MTRX_4_4 DOF2PD_MTRX_4_5
# MODULES DOF2PD_MTRX_4_6 DOF2PD_MTRX_4_7 DOF2PD_MTRX_4_8 DOF2PD_MTRX_4_9
# MODULES DOF2PD_MTRX_4_10 DOF2PD_MTRX_4_11 DOF2PD_MTRX_4_12 DOF2PD_MTRX_4_13
# MODULES DOF2PD_MTRX_4_14  
# MODULES 

  3054   Tue Jun 8 00:38:22 2010 Koji, KiwamuUpdateIOOimproved Gaussian beam in new IOO

The shape of the beam spot in the new input optics got much much better 

As Alberto and Kiwamu found on the last week, the beam spot after MMT1 had not been good. So far we postponed the mode measurement due to this bad beam profile.

Today after we did several things in the vacuum chamber, the beam spot became really a good Gaussian spot. See the attachment below.

There were two problems which had caused the bad profile:

(1)  a steering mirror after MMT1 with the incident angle of non 45 deg

(2) clipping at the Faraday.

 

Also MCT_QPD and MCT_CCD were recovered from misalignment  

Tomorrow we are going to restart the mode matching. 

 


(what we did)

* We started from checking the shape of the beam going out from the BS chamber. There still were some stripes which looked like an interference on the spot. 

* We found a steering mirror after MMT1 had the incident angle of non 45 deg. In fact the mirror had a large transmission. After we made the angle roughly 45 deg, the stripes disappeared.

However the spot still didn't look a good Gaussian, it looked slightly having a bump on the horizontal profile.

* Prior to moving of some optics in the vacuum, we ran the A2L_MC scripts in order to check the beam axis. And it was okay.

* To recover the MCT, we steered one of the vacuum mirrors which was located after the pick off mirror.  And after aligning some optics on the AP table, finally we got MCT recovered.

 * We rearranged MC_refl mirrors according to the new optical layout that Koji has made. At the same time the mirrors for IFO_refl was also rearranged coarsely.

 * We leveled the optical table of the MC chamber by moving some weights. Then we locked the MC again and aligned it. We again confirmed that the beam axis was still fine by running the A2L scripts.

 * We found the beam going through Faraday was off-centered by ~5mm toward the west. So we moved it so that the beam propagates on the center of it. 

 * Then looking at the beam profile after MMT1, we found that the profile became really nicer. It showed a beautiful Gaussian. 

In the attachment below, the top panel represents the horizontal profile and the bottom one represents the vertical profile.

The blue curves overlaid on the plot are fitted Gaussian profile, showing beautiful agreements with the measured profile.

  3053   Mon Jun 7 07:39:38 2010 AlbertoOmnistructureElectronicsCapacitor Bridge Test

Quote:

To get a feel for the Capacitive Bridge problems, we setup a simple bridge using fixed (1 nF) caps on a breadboard. We used an SR830 Lock-In amplifier to drive it and readout the noise.

The measurement setup for the Capacitor Bridge Test is still sitting on one of the work benches.

Unless the experiment is supposed to continue today, the equipment shouldn't have been left on the bench. It should have been  taken back to the lab.

Also the cart with HP network analyzer used for the test was left in the desk area. That shouldn't have left floating around in the desk area anyway.

The people responsible for that, are kindly invited to clean up after themselves.

  3052   Sun Jun 6 08:08:05 2010 rana, sanjitSummaryElectronicsCapacitor Bridge Test

To get a feel for the Capacitive Bridge problems, we setup a simple bridge using fixed (1 nF) caps on a breadboard. We used an SR830 Lock-In amplifier to drive it and readout the noise.

CapacitanceBridge.png

We measured the cap values with an LCR meter. They were all within a few % of 0.99 nF.

With a 0.5 V drive to the top of the bridge, the A-B voltage was ~2 mV as expected from the matching of the capacitors.

(** Note about the gain in the SR830: In order to find the magnitude of the input referred signal, one has to divide by G. G = (10 V)/ Sensitivity. 'Sensitivity' is the setting on the front panel.)

  1. Directly measuring from Vs to ground gives 0.5 V, as expected. This is done to verify the calibration later on.
  2. Shorting the A and B wires to ground gives ~0 V and lets us measure the noise. On the spectrum analyzer it was ~400 nV/rHz at 100 Hz and rising slowly to 4 uV/rHz at 100 mHz. In this state, the sensitivity was 10 mV, so the overall gain was 1000. That gives an input referred level of ~0.4 nV/rHz at the input.
  3. Hooking up now to A-B: the signal is ~10x larger than the 'dark' noise everywhere. 2 uV/rHz @ 100 Hz, 10 uV/rHz @ 10 Hz, 50 uV/rHz @ 1 Hz. The spectrum is very non-stationary; changing by factors of several up and down between averages. Probably a problem with the cheapo contacts in the breadboard + wind. The gain in this state was still 1000. So at 1 Hz, its 50 nV/rHz referred to the input.

To convert into units of capacitance fluctuation, we multiply by the capacitance of the capacitors (1 nF) and divide out by the peak-peak voltage (1 V). So the bridge sensitivity is 50e-9 * 1e-9 = 5 x 10^-17 F/rHz.

If we assume that we will have a capacitive displacement transducer giving 1 nF capacitance change for a 0.1 mm displacement, this bridge would have a sensitivity of 5 x 10^-12 m/rHz @ 1 Hz. We would like to do ~50-100x better than this. The next steps should be:

  1. Solder it all together on a PCB to have less air current sensitivity and decent contacts.
  2. Use a low-noise FET input. Since the impedance of the bridge is ~5 kOhms at this frequency, we are probably current noise limited.
  3. Estimate the oscillator amplitude noise sensitivity.
  3051   Sun Jun 6 04:48:41 2010 ranaUpdateCOCITM01 HR Phase Map

While trying to set up the SIS-FFT to use our new ITM phase maps, I noticed that the surface of our ITMs looks pretty good actually (even compared to the aLIGO pathfinder optic map on the AIC wiki). I'm attaching it here for your viewing pleasure.

The code to plot it has been added to the SVN in the PhaseMaps/mat directory.

  3050   Fri Jun 4 23:52:57 2010 ranaUpdatePEMseismometers off of linoleum floor

Untitled.pnghuddlez.png

For the huddle test, I have updated the code to divide the residual by sqrt(2) because of the assumption of equal noise from the 2 Guralps. We would have to multiply this trace by sqrt(2) to compare with the previous results.

Now the question is, how do I add a low noise ~50 mV offset to the front of the Guralp breakout box to test for the noise of the box?

  3049   Fri Jun 4 11:32:51 2010 Alberto, kiwamuUpdateIOOMC MMT1 Mirror Tests
[Alberto, Kiwamu]
Last Wednesday, we measured the beam profile after the MC mode matching telescope n.1 (MMT1). We found that the reflected beam had an irregular profile when observed with the beam scan. Fringes also appeared on an IR card.
We thought that such effect could be due to interference of the main reflected beam with the beam reflected by the back surface of the mirror.
 
To test the hypothesis we checked the transmitted and the reflected beams of a spare optic identical to MMT1. (This was the same optic that got dropped during the cleaning/baking process.)
 
We tested it on the PSL table, using a 200mW beam coming from the new 2W Innolight  laser. To maximize the separation between the two beams, we tested MMT1 at 45 degrees. The setup we used is shown here:
 
MCMMT1spareOpticsTestSetup.png
 
We looked at the beam reflected by MMT1 about 5 meters from the mirror. At that distance the beam spot had a size of about 1-2cm. it didn't look perfectly round, but it showed no fringes, as it had happened with original MMT1 inside the MC chamber.
At the transmission, the second ghost beam due to the back surface reflection (see picture above) was very week. In order to be able to see it on an IR card, we had to increase the laser pumping current from 1A to about 1.5A.
 
We are now thinking of a way to measure the relative power between the two. The problem is that they run very close to each other and it's not easy to resolve them with a power meter or a photodiode.
  3048   Thu Jun 3 22:33:31 2010 valeraSummaryCDSsimulated plant work

 I put matlab files and a summary into the 40m wiki for the fitting of the 40m Optickle transfer functions and generating digital filters for the simulated plant:

http://lhocds.ligo-wa.caltech.edu:8000/40m/Generating_DOF-%3EPD_digital_filters_based_on_Optickle_modeling

The filters are not loaded yet. Joe and Alex will make a rcg code to make a matrix of filters (currently 5x15=75 elements) which will enable the simulated plant tf's.

Joe and I tried to put a signal through the DARM loop but the signal was not going through the memory location in the scx part of the simulated plant.

 

Edit by Joe:

I was able to track it down to the spx model not running properly.  It needed the Burt Restore flag set to 1.  I hadn't done that since the last rebuild, so it wasn't actually calculating anything until I flipped that flag.  The data is now circulating all the way around.  If I turn on the final input (the same one with the initial 1.0 offset), the data circulates completely around and starts integrating up.  So the loop has been closed, just without all the correct filters in.

  3047   Thu Jun 3 22:17:05 2010 ranaUpdatePEMseismometers off of linoleum floor

 At ~2350 UTC on June 2, the seismometers were turned off. After the granite slab was repositioned with the new lead, the Ranger was turned on, but not the Guralps.

Now, after ~24 hours, I have put the Guralps onto the granite and turned them on. During this off time, the input channels should be ADC noise limited (or perhaps limited by the INA134 differential receiver chips inside of the Sander Liu AA chassis). The following plot shows that this noise level is ~0.8 uV/rHz and then rising like ~1/sqrt(f) below 5 Hz:

Untitled.png

I checked the slab again by whacking it. It still rings with a Q of several, so I think we need to make the lead flatter. There should barely be any room between the granite and the linoleum.

UPDATE:

I guessed that it should be possible to make the slab-to-floor coupling better with flatter lead (Brian Lantz suggested to use lead sheets). So I removed my booties and jumped up and down on the granite several times. Because of my soft sole shoes, I was able to make an impulsive impact without shattering the granite. The effect of the stomping was pretty dramatic - the horizontal resonance frequency has gone up by a factor of 2. The red trace shows the new TF after the stomping:

a.png

 And the resulting spectrum is here too. As you can see, there is no excess between the Ranger and the Guralps until ~50 Hz where the mechanical resonance in the short direction (non-MC dir) takes over.

Untitled.png

 So, the lesson for next time is to flatten the balls a little more. I leave it to Nancy to calculate the horizontal resonant frequencies of this lead/granite combo to see if it matches with our measurements.

  3046   Thu Jun 3 14:40:28 2010 Alberto, KiwamuUpdateIOOmode measurement of new input optics

Quote:

inside_vac_2.png

 

For the record, we wanted to check whether the fringes on the beam spot were caused by SM2 (see diagram above). We tried two different mirrors for SM2,

The first was one of the flat, 45 degree ones that were already on the BS table. The last, which is the one currently in place, was inside the plastic box with the clean optics that Jenne left us .

The fringes were present in both cases.

  3045   Thu Jun 3 14:13:20 2010 JenneUpdateIOOmode measurement of new input optics

Quote:

I checked the measured data of the mode profile which was taken on the last Tuesday.

For the vertical profile,

the plot shows a good agreement between the expected radius which is computed from the past measurement, and that measured on the last Tuesday.

However for the horizontal profile,

it looks like being overestimated. This disagreement may come from the interference imposed on the Gaussian spot as we've been worried. 

So we should solve this issue before restarting this mode matching work.

 - The next step we should do are;

checking the effect of MMT1 on the shape of the beam spot by using spare MMT1

 

NOTE:

The expected curve in the attached figure were computed by using the fitted parameter listed on the entry 2984 .

In the calculation the MMT1 is placed at 1911mm away from MC3 as we measured.

And the focal length of MMT1 is set to be f=-2500mm.

 When / if you use the other MMT1 mirror, make sure to take note whether or not it says "SPARE" on it in pencil.  I don't remember if it's the other MMT1, or if it's one of the MMT2's that says this.  The mirror was baked, so it's okay to use in the vacuum, however it's the one which was dropped on the floor (just prior to baking), so any discrepancies measured using that optic may not be useful.  I don't know how strong the CVI coatings are to scratches resulting in being dropped from a ~1m height.  Bob and I didn't see any obvious big scratches that day, but that doesn't necessarily give it a clean bill of health. 

The optic labeled "SPARE" should NOT be used as the final one in the IFO.

  3044   Thu Jun 3 14:01:51 2010 kiwamuUpdateIOOmode measurement of new input optics

I checked the measured data of the mode profile which was taken on the last Tuesday.

For the vertical profile,

the plot shows a good agreement between the expected radius which is computed from the past measurement, and that measured on the last Tuesday.

However for the horizontal profile,

it looks like being overestimated. This disagreement may come from the interference imposed on the Gaussian spot as we've been worried. 

So I guess we should solve this issue before restarting this mode matching work.

 - The next step we should do are;

checking the effect of MMT1 on the shape of the beam spot by using a spare of MMT1

 

NOTE:

The expected curve in the attached figure were computed by using the fitted parameter listed on the entry 2984 .

In the calculation the MMT1 is placed at 1911mm away from MC3 as we measured.

And the focal length of MMT1 is set to be f=-2500mm.

  3043   Thu Jun 3 13:14:27 2010 kiwamuUpdateIOOmode measurement of new input optics

I corrected the sketch of the new IOO. 

The sketch in the last entry was also replaced by the new one. 

http://nodus.ligo.caltech.edu:8080/40m/3029 

Quote:

Just note that MMT1 has RoC of -5m (negative!). This means that it is a lens with f=-2.5 m,

  3042   Thu Jun 3 00:47:17 2010 KevinUpdatePSL2W Beam Profile of Second Reflected Beam

[Koji, Kevin]

The profile of the Innolight 2W was previously measured by measuring the reflected beam from the front surface of a W2 window (see entry). To investigate thermal effects, Rana suggested also measuring the profile of the beam reflected from the back surface of the W2.

I used the same setup and methods as were used in the first measurement. The mirror was moved so that only the beam reflected from the back surface of the W2 was reflected from the mirror. This beam was reflected from both the front of the mirror and the back of the mirror. An extra beam dump was positioned to block the reflection from the back of the mirror.

This measurement was made with 2.004 A injection current and 25.04°C laser crystal temperature.

The data was fit to w = sqrt(w0^2+lambda^2*(x-x0)^2/(pi*w0)^2) with the following results

For the horizontal beam profile:

reduced chi^2 = 5.1

x0 = (-186 ± 6) mm

w0 = (125.8 ± 1.4) µm


For the vertical beam profile:

reduced chi^2 = 14.4

x0 = (-202 ± 11) mm

w0 = (132.5 ± 2.7) µm


In the following plots, the blue curve is the fit to the vertical beam radius, the purple curve is the fit to the horizontal beam radius, * denotes a data point from the vertical data, and + denotes a data point from the horizontal data.

  3041   Wed Jun 2 22:58:04 2010 KevinUpdatePSL2W Second Reflected Beam Profile

[Koji, Kevin]

The profile of the Innolight 2W was previously measured by measuring the reflected beam from the front surface of a W2 window (see entry). To investigate thermal effects, Rana suggested also measuring the profile of the beam reflected from the back surface of the W2.

I used the same setup and methods as were used in the first measurement. The mirror was moved so that only the beam reflected from the back surface of the W2 was reflected from the mirror. This beam was reflected from both the front of the mirror and the back of the mirror. An extra beam dump was positioned to block the reflection from the back of the mirror.

This measurement was made with 2.004 A injection current and 25.04°C laser crystal temperature.

The data was fit to w = sqrt(w0^2+lambda^2*(x-x0)^2/(pi*w0)^2) with the following results

For the horizontal beam profile:

reduced chi^2 = 5.1

x0 = (-186 ± 6) mm

w0 = (125.8 ± 1.4) µm

For the vertical beam profile:

reduced chi^2 = 14.4

x0 = (-202 ± 11) mm

w0 = (132.5 ± 2.7) µm

In the following plots, the blue curve is the fit to the vertical beam radius, the purple curve is the fit to the horizontal beam radius, * denotes a data point from the vertical data, and + denotes a data point from the horizontal data.

The differences between the beam radii for the beam reflected from the front surface and the beam reflected from the back surface are

Δw0_horizontal = (12.8 ± 1.6) µm

Δw0_vertical = (8.5 ± 2.9) µm

So the two measurements are not consistent. This suggests that the passage through the W2 altered the profile of the beam.

  3040   Wed Jun 2 22:25:39 2010 KevinUpdatePSLLow Power 2W Beam Profile

Koji is worried about thermal lensing introducing errors to the measurement of the 2W beam profile so I measured the profile at a lower power.

I used the same setup and methods used to measure the profile at 2W (see entry). This measurement was taken with an injection current of 1.202 A and a laser crystal temperature of 25.05° C. This corresponds to approximately 600 mW (see power measurement).

The data was fit to  w = sqrt(w0^2+lambda^2*(x-x0)^2/(pi*w0)^2) with the following results

For the horizontal beam profile:

reduced chi^2 = 2.7

x0 = (-203 ± 3) mm

w0 = (151.3 ± 1.0) µm

For the vertical beam profile:

reduced chi^2 = 6.8

x0 = (-223 ± 6) mm

w0 = (167.5 ± 2.2) µm

In the following plots, the blue curve is the fit to the vertical beam radius, the purple curve is the fit to the horizontal beam radius, * denotes a data point from the vertical data, and + denotes a data point from the horizontal data.

The differences between the beam radii for the low power and high power measurements are

Δw0_horizontal = (38.3 ± 1.2) µm

Δw0_vertical = (43.5 ± 2.4) µm

Thus, the two measurements are not consistent. To determine if the thermal lensing is in the laser itself or due to reflection from the W2 and mirror, we should measure the beam profile again at 2W with a razor blade just before the W2 and a photodiode to measure the intensity of the reflection off of the front surface. If this measurement is consistent with the measurement made with the beam scan, this would suggest that the thermal lensing is in the laser itself and that there are no effects due to reflection from the W2 and mirror. If the measurement is not consistent, we should do the same measurement at low power to compare with the measurement described in this entry.


  3039   Wed Jun 2 21:21:43 2010 steveUpdatePEMseismometers off of linoleum floor

Quote:

Steve for Nancy,

Seismometer interface box ac power was turned off, Guralps disconnected and moved. Ranger locked, moved and released. Nancy will describe the rest soon.

The flattened lead balls were checked for their heights by the calliper, and were all in the range of 9.50 to 9.70 mm.

The rechecking was done by using these balls between two aluminium plates and checking their levelling. When confirmed this, we proceeded to install the balls(no more balls :P ) in their place.

The Guralps were switched off by switching off the power supply to them. The ranger mass was clamped in order to be able to move it. This can be undone by rotating the transport screw counter-clockwise.

We installed the flattened lead ballsin the space made for them. The granite was then placed on it with the help of many other people in the lab.  It was lowered by hanging it on two straps held by people , and then placed in the space marked for it.

Did we then turn on the seismometers? Did we release the locking screw on the Ranger? What happened to Bat-Boy??? Since they make a good mystery I will choose to leave them out of my elog entry.

  3038   Wed Jun 2 18:36:20 2010 valeraDAQCDSNoise generators in LSP

Alex wrote a new code to implement LSP noise generator. The code is based on 64 bit random number generator from Numerical Recipes 3rd ed ch 7.1 (p 343).

Joe made two instances in the LSP model.

The attached plot shows the spectra and coherence of two generators. The incoherence is ~1/Navg - statistically consistent with no coherence.

  3037   Wed Jun 2 18:09:32 2010 steveUpdatePEMseismometers off of linoleum floor

Steve for Nancy,

Seismometer interface box ac power was turned off, Guralps disconnected and moved. Ranger locked, moved and released. Nancy will describe the rest soon.

 

  3036   Wed Jun 2 17:34:33 2010 josephb, alex, valeraUpdateCDSCDS updates

From what I understand, Alex rewrote portions of the framebuilder and testpoint codes and then recompiled them in order to get more than 1 testpoint per front end working.   I've tested up to 5 testpoints at once so far, and it worked.

We also have a new noise component added to the RCG code.  This piece of code uses the random number generator from chapter 7.1 of Numerical Recipies Third Edition to generate uniform numbers from 0 to 1.  By placing a filter bank after it should give us sufficient flexibility in generating the necessary noise types.  We did a coherence test between two instances of this noise piece, and they looked pretty incoherent.  Valera will add a picture of it when it finishe 1000 averages to this elog.

I'm in the process of propagating the old suspension control filters to the new RCG filter banks to give us a starting point.  Tomorrow Valera and I are planning to choose a subset of the plant filters  and put them in, and then work out some initial control filters to correspond to the plant.  I also need to think about adding the anti-aliasing filters and whitening/dewhitening filters.

 

  3035   Wed Jun 2 11:28:31 2010 KojiUpdateMOPAlaser headtemp is up

Last night we stopped the air conditioning. It made HDTEMP increase.
Later we restored them and the temperature slowly recovered. I don't know why the recovery was so slow.

Quote:

Is the cooling line clogged? The chiller temp is 21C See 1 and 20 days plots

 

  3034   Wed Jun 2 11:25:16 2010 josephb,alexUpdateCDSCDS saga (aka the bad code saga)

Alex updated the awg.par file to handle all the testpoints.  Basically its very similar to the testpoint.par, but the prognum lines have to be 1 higher than the corresponding prognum in testpoint.par.  A entry looks like:

[C1-awg0]
hostname=192.168.1.2
prognum=0x31001002

After running "diag -i" and seeing some RPC number conflicts, we went into /cvs/cds/caltech/cds/target/gds/param/diag_C.conf and changed the line from

&chn * *  192.168.1.2 822087685 1

to

&chn * *  192.168.1.2 822087700 1

The number represents an RPC number.  This was conflicting with the RPC number associated with the awgtpman processes.  We then had to update the /etc/rpc file as well.  At the end we changed chnconf 822087685 to chnconf 822087700.  We then run /usr/sbin/xinetd reload

Lastly we edited the /etc/xinetd.d/chnconf file line

server_args             = /cvs/cds/caltech/target/gds/param/tpchn_C4.par /cvs/cds/caltech/target/gds/param/tpchn_C5.par

to

server_args             = /cvs/cds/caltech/target/gds/param/tpchn_C1.par /cvs/cds/caltech/target/gds/param/tpchn_C2.par /cvs/cds/caltech/target/gds/param/tpchn_C3.par /cvs/cds/caltech/target/gds/param/tpchn_C4.par /cvs/cds/caltech/target/gds/param/tpchn_C5.par /cvs/cds/caltech/target/gds/param/tpchn_C6.par /cvs/cds/caltech/target/gds/param/tpchn_C7.par /cvs/cds/caltech/target/gds/param/tpchn_C8.par /cvs/cds/caltech/target/gds/param/tpchn_C9.par

 

Alex also recompiled the frame builder code to be able to handle more than 7 front ends.  This involved tracking down a newer version of libtestpoint.so on c1iscex and moving it over to megatron, then going in and by hand adding the ability to have up to 10 front ends connected.

Alex has said he doesn't like this code and would like it to dynamically allocate properly for any number of servers rather than having a dumb hard coded limit.

Other changes he needs to make:

1) Get rid of set dcu_rate ## = 16384 type lines in the daqrc file.  That information is available from the /caltech/chans/C1LSC.ini type files which are automatically generated when you compile a model.  This means not having to go in by hand to update these in daqrc.

2) Get some awg.par and testpoint.par rules, so that these are automatically updates when you build a model.  Make it so it automatically assigns a prognum when read in rather than having to hard code them in by hand.

3)Slave the awgtpmans to a single clock running from the IO processor x00. This ensures they are all in sync.

 

 

 

  3033   Wed Jun 2 07:54:55 2010 steveUpdateMOPAlaser headtemp is up

Is the cooling line clogged? The chiller temp is 21C See 1 and 20 days plots

  3032   Wed Jun 2 04:27:02 2010 KojiUpdatePSL2W Beam Profile

This is what I already told to Kevin and Rana:

A direct output beam is one of the most difficult measurements for the mode profiling.
I worried about the thermal lensing.

Since most of the laser power goes through the substrate (BK7) of the W2 window, it may induce thermal deformation on the mirror surface.
An UV fused silica window may save the effect as the thermal expansion coefficient is 0.55e-6/K while BK7 has 7.5e-6.

In addition to the thermal deformation issue, the pick-off setup disables us to measure the beam widths near the laser aperture.

I rather prefer to persist on the razor blade then use the pick off between the blade and the PD.

I also confess that the description above came only from my knowledge, and not from any scientific confirmation including any calculation.
If we can confirm the evidence (or no evidence) of the lensing, it is a great addition to my experience.

Quote:

[Rana, Kiwamu, Kevin]

The Innolight 2W beam profile was measured with the beam scan. A W2-IF-1025-C-1064-45P window was used to reflect a small amount of the main beam. A 5101 VIS mirror was used to direct just the beam reflected from the front surface of the W2 down the table (the beam reflected from the back surface of the W2 hit the optic mount for the mirror). A razor blade beam dump was used to stop the main transmitted beam from the W2. The distance from the laser was measured from the front black face of the laser to the front face of the beam scan (this distance is not the beam path length but was the easiest and most accurate distance to measure). The vertical and horizontal beam widths were measured at 13.5% of the maximum intensity (each measurement was averaged over 100 samples). These widths were divided by 2 to get the vertical and horizontal radii.

The mirror was tilted so that the beam was close to parallel to the table. (The center of the beam fell by approximately 2.1 mm over the 474 mm that the measurement was made in).

The measurement was taken with an injection current of 2.004 A and a laser crystal temperature of 25.04°C.

This data was fit to w = sqrt(w0^2+lambda^2*(x-x0)^2/(pi*w0)^2) with lambda = 1064nm with the following results

For the horizontal beam profile:

reduced chi^2 = 4.0

x0 = (-138 ± 3) mm

w0 = (113.0 ± 0.7) µm

For the vertical beam profile:

reduced chi^2 = 14.9

x0 = (-125 ± 4) mm

w0 = (124.0 ± 1.0) µm

In the following plots, the blue curve is the fit to the vertical beam radius, the purple curve is the fit to the horizontal beam radius, * denotes a data point from the vertical data, and + denotes a data point from the horizontal data.

 

ELOG V3.1.3-